OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x.

OpenStudy (anonymous):

x+x+y=46 2x+y=46--(*) y=4+x put y=4+x into (*) then we can get x Sub x into y=4+x, we will get the value of y